# Dream-IF: Dynamic Relative EnhAnceMent for Image Fusion

### v1

@@ 4.2 Comparison without Degradation
Quantitative Comparison. The quantitative evaluation, conducted using five widely recognized metrics across the LLVIP and MSRS datasets, is summarized in Table 1. On the LLVIP dataset, our method achieves the highest scores across all five metrics, namely EI, EN, AG, $Q^{AB/F}$, and VIFF. Notably, the highest EN and AG scores indicate that our method retains the richest information and sharpest details, attributed to its ability to enhance non-dominant regions while preserving dominant features dynamically. The superior $Q_{AB/F}$ score reflects better alignment of local gradients and intensities between the source and fused images, demonstrating our method’s capability to effectively integrate multi-modal information. Furthermore, the highest VIFF score confirms that the fused images retain a substantial amount of visual information from the source modalities, underscoring the high perceptual quality of our results. These quantitative findings align with the qualitative observations, validating that our method achieves superior fusion performance through dynamic relative enhancement and robust multi-modal integration. Our method also demonstrates competitive performance across most evaluation metrics on the MSRS dataset. The consistency across datasets further validates the generalizability and robustness of our approach under diverse scenarios.
@@
Table 2: Comparison with SOTAs under degradation. The red indicates the best.
Methods EI $\uparrow$ AG $\uparrow$ PSNR $\uparrow$ $Q^{AB/F}$$\uparrow$ VIFF $\uparrow$
STDFusion [19] $6.65$ $2.65$ $31.45$ $0.19$ $0.24$
TarDAL [17] $12.94$ $5.43$ $27.14$ $0.23$ $0.22$
CDDFuse [43] $10.89$ $4.23$ $30.95$ $0.34$ $0.39$
EMMA [44] $10.62$ $4.07$ $31.65$ $0.36$ $0.42$
Text-IF [34] $11.03$ $4.22$ $30.91$ $0.38$ $0.44$
TTD [1] $10.89$ $4.23$ $30.95$ $0.34$ $0.39$
Text-DiFuse  [36] $9.92$ $3.78$ $25.83$ $0.29$ $0.39$
Dream-IF(Ours) ${\color[rgb]{1,0,0}\mathbf{14.95}}$ ${\color[rgb]{1,0,0}\mathbf{5.87}}$ ${\color[rgb]{1,0,0}\mathbf{31.88}}$ ${\color[rgb]{1,0,0}\mathbf{0.64}}$ ${\color[rgb]{1,0,0}\mathbf{0.72}}$
4.3 Comparison with Degradation
To evaluate the robustness and practicality of our method in real-world scenarios, we introduce a diverse set of degradations [37] to simulate challenging conditions that amplify the disparity between modalities. These degradations include gaussian noise, poison noise, speckle noise, etc., which are commonly encountered in practical applications. To ensure fairness, we compare existing methods for image fusion and restoration. This includes two-stage fusion, which performs degradation recovery followed by fusion, and one-stage fusion with degradation.
@@ 4.3 Comparison with Degradation
Figure 5: Qualitative comparisons of object detection on the LLVIP dataset.
@@
Two-stage Fusion. For blind restoration, we apply random degradation to the LLVIP test set. To ensure a fair comparison, we employ a two-stage fusion strategy. Specifically, for all the competing methods, we adopt SCUNet [38], a widely recognized restoration model, to restore images before fusion. For comparison, our Dream-IF directly performs image restoration and fusion in a unified model, which is a more challenging.
@@
Quantitative results. Our method compared with the combined results of SOTA image fusion methods after SCUNet restoration on degraded source images are presented in Fig. 4. Our method retains more details, exhibiting a more realistic overall texture, particularly in the green box. CDDFuse exhibits competitive performance but also suffers from detail and texture loss. TarDAL suffers from fidelity loss, while Text-IF, TTD, and Text-DiFuse produce excessively smooth results, omitting important information. Although STDFusion and EMMA perform competitively in the green box, they fail to preserve details in the red box as effectively as ours. The advantages demonstrate that our approach effectively performs both restoration and fusion, with its strengths becoming more pronounced as the differences between modalities increase.
@@
Qualitative results are reported in Table 2, demonstrating that our method achieves competitive results even without the use of a specialized restoration model. Specifically, our method achieves optimal values across all metrics. The highest EI and AG scores indicate that Dream-IF excels at retaining texture and detail. VIFF and $Q^{AB/F}$ indicate that our method effectively preserves mutual information between the source images, ensuring details and contextual information are retained in the output while maintaining subjective visual quality. It underscores the ability of our approach to preserve image fidelity and enhance degraded features without relying on a dedicated restoration model.
@@
One-stage Fusion. To verify the performance of our method in simultaneous fusion and restoration within a single framework, we further evaluate our Dream-IF with Text-IF, a text-guided restoration and fusion model. Since Text-IF requires descriptive text to specify the degradation type, we provide degradation text guidance for Text-IF. It is worth noting that our model does not take any prior knowledge of the degradation type, which is a much more challenging setting to handle blind restoration and fusion. We evaluated image restoration using the learned perceptual image patch similarity (LPIPS) metric. LPIPS captures perceptual differences between images, with lower scores indicating greater similarity.
@@
Qualitative results. As presented in Fig. 6, it demonstrates that our method recovers more texture details affected by degradation. However, even with degradation descriptions, Text-IF fails to fully restore images affected by Gaussian noise, and it results in overly smooth fused images when applied to other types of noise.. Text-DiFuse removes noise through diffusion, but yield poor performance. This demonstrates that our method is robust to various degradations, benefiting from dynamic dominance enhancement to improve inferior features. In contrast, our method leverages information from both modalities to seamlessly integrate image fusion and restoration, enabling effective blind restoration and fusion. It further validates our effectiveness in utilizing the comprehensive nature of fusion to drive overall image enhancement in the multi-modal fusion task.
@@
Methods Triplet Gaussian Poisson Speckle
Text-IF* $0.445$ $0.440$ $0.281$ $0.270$
Text-DiFuse $0.442$ $0.439$ $0.263$ $0.261$
Dream-IF $\mathbf{0.288}$ $\mathbf{0.287}$ $\mathbf{0.252}$ $\mathbf{0.249}$
@@
Base $CE$ $SE$ EI $\uparrow$ AG $\uparrow$ PSNR $\uparrow$ $Q^{AB/F}$uparrow VIFF $\uparrow$
$\checkmark$ $13.82$ $5.48$ $31.93$ $0.66$ $0.72$
$\checkmark$ $\checkmark$ $14.32$ $5.71$ $31.68$ $0.71$ $0.83$
$\checkmark$ $\checkmark$ $14.52$ $5.81$ $31.59$ $0.72$ $0.85$
$\checkmark$ $\checkmark$ $\checkmark$ $\mathbf{15.17}$ $\mathbf{6.07}$ $\mathbf{32.33}$ $\mathbf{0.73}$ $\mathbf{0.86}$
4.4 Performance on Downstream Task
We perform experiments on object detection to verify the compatibility of Dream-IF with high-level downstream tasks. Specifically, we train the YOLOv5 [13] detector on the LLVIP dataset, using fusion results generated by competitive comparison methods, and evaluate performance using recall and mAP. As shown in Table 5, Dream-IF outperforms the competition in object detection, achieving the best performance in recall, mAP@.5, and mAP.5:.95. Additionally, Fig. 5 shows that most comparison methods can identify the main target but fail to detect smaller or incomplete targets. In contrast, our method successfully detects all objects in the scene, while other methods either miss detections or yield false positives, demonstrating Dream-IF’s enhanced detection accuracy, particularly for hard-to-detect targets.
@@ 4.4 Performance on Downstream Task
Methods Recall mAP@.5 mAP@.5:.95
STDFusion [19] $0.894$ $0.954$ $0.602$
CDDFuse  [43] $0.894$ $0.954$ $0.602$
EMMA [44] $0.905$ $0.947$ $0.634$
Text-IF [34] $0.900$ $0.964$ $0.622$
TTD [1] $0.869$ $0.929$ $0.574$
Text-DiFuse [36] $0.889$ $0.952$ $0.609$
Dream-IF (Ours) $\mathbf{0.914}$ $\mathbf{0.968}$ $\mathbf{0.643}$
@@
4.5 Ablation Study
We conduct ablation studies on the LLVIP dataset, as shown in Table 4, to validate the effectiveness of the proposed CE and SE module. Specifically, we evaluate the performance of our model by removing the CE and SE modules, treating the model without these components as the baseline. The results demonstrate that adding either the CE or SE module leads to minor improvements in the performance across all evaluation metrics. When both modules are incorporated into the model, substantial improvements in all indicators are observed. These results highlight the distinct and complementary contributions of each module. Notably, the inclusion of both CE and SE along with the final model configuration leads to the highest scores across all metrics, confirming that our method achieves the best qualitative and quantitative performance among all ablation settings. This further validates the importance of the CE and SE modules, as well as their combined effect in enhancing the overall model’s capabilities. The improvements across various indicators, clearly demonstrate the efficacy of our approach in achieving superior results.
5 Conclusion

### v2

@@ Introduction
Quantitative Comparison. The quantitative evaluation, conducted using five widely recognized metrics across the LLVIP, MSRS, and M3FD datasets, is summarized in Table 1. On the LLVIP dataset, our method achieves the highest scores across most metrics, namely AG, SF, $Q^{abf}$, and SSIM. Notably, the highest AG and SF scores indicate that our method retains the richest information and sharpest details, attributed to its ability to enhance non-dominant regions while preserving dominant features dynamically. The superior $Q_{abg}$ score reflects better alignment of local gradients and intensities between the source and fused images, demonstrating our method’s capability to effectively integrate multi-modal information. Furthermore, the highest SSIM score confirms that the fused images retain a substantial amount of visual information from the source modalities, underscoring the high perceptual quality of our results. These quantitative findings align with the qualitative observations, validating that our method achieves superior fusion performance through dynamic relative enhancement and robust multi-modal integration. Our method also demonstrates competitive performance across most evaluation metrics on the MSRS and M3FD datasets. The consistency across datasets further validates the generalizability and robustness of our approach under diverse scenarios.
@@
Methods AG SF $Q^{abf}$ SSIM VIFF
TarDAL $4.745$ $17.337$ $0.266$ $0.503$ $0.256$
CDDFuse $4.530$ $15.929$ $0.523$ $1.179$ $0.618$
EMMA $4.593$ $14.818$ $0.499$ $1.186$ $0.592$
Text-IF $4.863$ $16.554$ $0.536$ $1.156$ $0.609$
TTD $4.948$ $16.637$ $0.489$ $1.144$ $0.533$
Text-DiFuse $4.084$ $13.777$ $0.364$ $1.049$ $0.490$
FusionBooster $4.524$ $13.401$ $0.380$ $1.017$ $0.533$
Dream-IF $\mathbf{5.243}$ $\mathbf{17.441}$ $\mathbf{0.580}$ $\mathbf{1.186}$ $\mathbf{0.705}$
Table 2: Comparison with SOTAs under degradation. The bold indicates the best.
Comparison with Degradation
To assess the robustness and real-world applicability of our method, we introduce a diverse set of degradations (Zhang et al. 2021) to simulate challenging scenarios that amplify modality discrepancies. These degradations include Gaussian noise, Poisson noise, speckle noise, and other common artifacts encountered in practical applications (e.g., blur, etc.). Additionally, we incorporate synthetic rainy degradations to further mimic real-world conditions. For a fair comparison, we compare existing methods for image fusion and restoration. This includes two-stage fusion, which performs degradation recovery followed by fusion, and one-stage fusion with degradation.
@@
Two-stage Fusion. For blind restoration, we apply random degradation to the LLVIP test set. To ensure a fair comparison, we employ a two-stage fusion strategy. Specifically, for all the competing methods, we adopt SCUNet (Zhang et al. 2023), a widely recognized restoration model, to restore images before fusion. For comparison, our Dream-IF directly performs image restoration and fusion in a unified model, which is a more challenging.
@@
Quantitative results. Our method compared with the combined results of SOTA image fusion methods after SCUNet restoration on degraded source images are presented in Fig. 4. Our method retains more details, exhibiting a more realistic overall texture, particularly in the green box. CDDFuse exhibits competitive performance but also suffers from detail and texture loss. TarDAL suffers from fidelity loss, while Text-IF, TTD, Text-DiFuse and FusionBooster produce excessively smooth results, omitting important information. Although EMMA performs competitively in the green box, it fails to preserve details in the red box as effectively as ours. The advantages demonstrate that our approach effectively performs both restoration and fusion, with its strengths becoming more pronounced as the differences between modalities increase.
@@
Qualitative results are reported in Table 2, demonstrating that our method achieves competitive results even without the use of a specialized restoration model. Specifically, our method achieves optimal values across all metrics. The highest AG and SF scores indicate that Dream-IF excels at retaining texture and detail. VIFF, SSIM, and $Q^{abf}$ indicate that our method effectively preserves mutual information between the source images, ensuring details and contextual information are retained in the output while maintaining subjective visual quality. It underscores the ability of our approach to preserve image fidelity and enhance degraded features without relying on a dedicated restoration model.
@@
One-stage Fusion. To verify the performance of our method in simultaneous fusion and restoration within a single framework, we further evaluate our Dream-IF with Text-IF, a text-guided restoration and fusion model. Since Text-IF requires descriptive text to specify the degradation type, we provide degradation text guidance for Text-IF. It is worth noting that our model does not take any prior knowledge of the degradation type, which is a much more challenging setting to handle blind restoration and fusion. We evaluated image restoration using the learned perceptual image patch similarity (LPIPS). LPIPS captures perceptual differences between images, with lower scores indicating greater similarity.
Methods Rain Gaussian Poisson Speckle Blur Resize
Text-IF* $0.273$ $0.275$ $0.275$ $0.270$ $0.269$ $0.244$
Text-DiFuse $0.296$ $0.291$ $0.268$ $0.266$ $0.290$ $0.268$
Dream-IF $\mathbf{0.241}$ $\mathbf{0.255}$ $\mathbf{0.242}$ $\mathbf{0.240}$ $\mathbf{0.260}$ $\mathbf{0.241}$
@@
Qualitative results. As presented in Fig. 5, it demonstrates that our method recovers more texture details affected by degradation. However, even with degradation descriptions, Text-IF fails to restore images affected by Gaussian noise fully, and it results in overly smooth fused images when applied to other types of noise. Text-DiFuse removes noise through diffusion, but yields poor performance. This demonstrates that our method is robust to various degradations, benefiting from dynamic dominance enhancement to improve inferior features. In contrast, our method leverages information from both modalities to seamlessly integrate image fusion and restoration, enabling effective blind restoration and fusion. It further validates our effectiveness in utilizing the comprehensive nature of fusion to drive overall image enhancement in the multi-modal fusion task.
@@
Base RD CE SE EI AG PSNR $Q^{abf}$ VIFF
$\checkmark$ - - - $5.402$ $18.032$ $0.627$ $1.181$ $0.736$
$\checkmark$ $\checkmark$ - - $5.638$ $18.594$ $0.659$ $1.185$ $0.756$
$\checkmark$ $\checkmark$ $\checkmark$ - $5.642$ $18.633$ $0.675$ $1.193$ $0.770$
$\checkmark$ $\checkmark$ - $\checkmark$ $5.746$ $19.446$ $0.668$ $1.191$ $0.797$
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\mathbf{5.926}$ $\mathbf{19.992}$ $\mathbf{0.679}$ $\mathbf{1.198}$ $\mathbf{0.800}$
@@
Methods Recall mAP@.5 mAP@.5:.95
TarDAL $0.8850$ $0.9495$ $0.6291$
CDDFuse $0.8927$ $0.9496$ $0.6031$
EMMA $0.8894$ $0.9502$ $0.6192$
Text-IF $0.8951$ $0.9517$ $0.6148$
TTD $0.8888$ $0.9497$ $0.6252$
Text-DiFuse $0.8742$ $0.9394$ $0.5964$
FusionBooster $0.8925$ $0.9502$ $0.6245$
Dream-IF $\mathbf{0.9041}$ $\mathbf{0.9543}$ $\mathbf{0.6327}$
@@
Performance on Downstream Task
We perform experiments on object detection to verify the compatibility of Dream-IF with high-level downstream tasks. Specifically, we train the YOLOv11 (Khanam and Hussain 2024) detector on the LLVIP dataset, using fusion results generated by competitive comparison methods, and evaluate performance using recall and mAP. As shown in Table 5, Dream-IF outperforms the competition in object detection, achieving the best performance in recall, mAP@.5, and mAP.5:.95. Additionally, visualizations of the comparisons in object detection are shown in the Appendix.
Ablation Study
We conduct ablation studies on the LLVIP dataset, as shown in Table 4, to validate the effectiveness of the proposed CE and SE module. Specifically, we evaluate the performance of our model by removing the CE and SE modules, treating the model without these components as the baseline. The results demonstrate that adding either the CE or SE module leads to minor improvements in the performance across all evaluation metrics. When both modules are incorporated into the model, substantial improvements in all indicators are observed. These results highlight the distinct and complementary contributions of each module. Notably, the inclusion of both CE and SE along with the final model configuration leads to the highest scores across all metrics, confirming that our method achieves the best qualitative and quantitative performance among all ablation settings. This further validates the importance of the CE and SE modules, as well as their combined effect in enhancing the overall model’s capabilities. The improvements across various indicators demonstrate the efficacy of our approach in achieving superior results.
Conclusion
@@
Acknowledgments
This work was sponsored by the National Natural Science Foundation of China (No.s 62222608, 62436002, 62476198), the Tianjin Natural Science Funds for Distinguished YoungScholar (No. 23JCJQJC00270), the Zhejiang ProvincialNatural Science Foundation of China (No. LD24F020004), and the Natural Science Foundation of Tianjin (No. 25JCYBJC00950).